# Randomized Black-Box PIT for Small Depth +-Regular Non-commutative Circuits

### v1

@@ C.1 Substitution Automaton for Commutative Transformation
    It is important to note that $f_{\lambda}\neq 0$. While $\tilde{f}\neq f_{\lambda}$, all monomials of $A^{\tilde{m}}_{{\hat{f}}}$ are present in $f_{\lambda}$ with their coefficients scaled by exactly $D_{2}$. However, at least one monomial $m_{l}\in B_{F}^{\tilde{m}}$ has its coefficient in $\tilde{f}$ scaled by at most $(D_{2}-1)$.
    Since only monomials in $A^{\tilde{m}}_{{\hat{f}}}\sqcup B_{F}^{\tilde{m}}$ are transformed into the commutative monomial $\tilde{m}$ after product sparsification and commutative transformation (Steps 2 and 3), performing these steps on $f_{\lambda}$, will yield a non-zero commutative polynomial. In particular, the coefficient of $\tilde{m}$ is non-zero.
    As only monomials in $A^{\tilde{m}}_{{\hat{f}}}\sqcup B_{F}^{\tilde{m}}$ are transformed into the commutative monomial $\tilde{m}$ after product sparsification and commutative transformation (Steps 2 and 3), if we carry out product sparsification and commutative transformation on $f_{\lambda}$, the resulting commutative polynomial is non-zero. In particular, the coefficient of $\tilde{m}$ is non-zero in the transformed commutative polynomial.
@@ D.2 Substitution Automaton for Case 2
Given that each monomial of $\tilde{f}$ is a product of $\xi$-patterns (see Defintion 4.0.2), it is logical to explain the automaton’s guesses in terms of these $\xi$-patterns.
@@
Here, $c_{j,\rho}$ and $c_{j,I}$ are commutative monomials over $Y\sqcup Z$ defined by the path $\rho$ and the set $I$, respectively. The last two equalities hold because, $I_{\rho,j}$ can be any subset of size $s-1$ and is independent of $I_{\rho,k}$ for $k\neq j$ (where $k\leq D_{2}$).
@@
The last equality holds because if $m\in Mon(f)$, it can be written as $m=m_{1}\cdot m_{2}\cdots m_{D_{2}}$ and for some $i\in[s]$ such that each sub-monomial $m_{j}\in X^{D_{1}}$, has a non-zero coefficient in $Q_{ij}$. Also suppose $m$ is computed by $\prod_{j\in[D_{2}]}Q_{i,j}$ for some $i\in[s]$ then by definition $m\in Mon(f)$. This completes the proof. ∎

### v2

@@ C.1 Substitution Automaton for Commutative Transformation
    It is important to note that $f_{\lambda}\neq 0$. While $\tilde{f}\neq f_{\lambda}$, all monomials of $A^{\tilde{m}}_{{\hat{f}_{1}}}$ are present in $f_{\lambda}$ with their coefficients scaled by exactly $D_{2}$. However, at least one monomial $m_{l}\in B_{F_{1}}^{\tilde{m}}$ has its coefficient in $\tilde{f}$ scaled by at most $(D_{2}-1)$.
    Since only monomials in $A^{\tilde{m}}_{{\hat{f}_{1}}}\sqcup B_{F_{1}}^{\tilde{m}}$ are transformed into the commutative monomial $\tilde{m}$ after product sparsification and commutative transformation (Steps 2 and 3), performing these steps on $f_{\lambda}$, will yield a non-zero commutative polynomial. In particular, the coefficient of $\tilde{m}$ is non-zero.
    As only monomials in $A^{\tilde{m}}_{{\hat{f}_{1}}}\sqcup B_{F_{1}}^{\tilde{m}}$ are transformed into the commutative monomial $\tilde{m}$ after product sparsification and commutative transformation (Steps 2 and 3), if we carry out product sparsification and commutative transformation on $f_{\lambda}$, the resulting commutative polynomial is non-zero. In particular, the coefficient of $\tilde{m}$ is non-zero in the transformed commutative polynomial.
@@ D.2 Substitution Automaton for Case 2
Given that each monomial of $\tilde{f}$ is a product of $\xi$-patterns (see Definition 4.0.2), it is logical to explain the automaton’s guesses in terms of these $\xi$-patterns.
@@
$=$ $\prod_{j=1}^{D_{2}-1}\left(\sum\limits_{\rho:q_{0}\overset{m_{j}}{\leadsto}q_{0}}c_{j,\rho}\cdot\xi_{I_{\rho,j}}\right)\times\left(\sum\limits_{\rho:q_{0}\overset{m_{D_{2}}}{\leadsto}q_{s-1}}c_{j,\rho}\cdot\xi_{I_{\rho,j}}\right)$
@@
The last two equalities hold because, $I_{\rho,j}$ can be any subset of size $s-1$ and is independent of $I_{\rho,k}$ for $k\neq j$ (where $k\leq D_{2}$).
@@
Let $Q_{i,j,I}$ be the set of all monomials in $\left(\sum_{\begin{subarray}{c}I\subseteq[D_{1}],|I|=s-1\end{subarray}}c_{j,I}\xi_{I}\right)$ that has the same non-commutative part $\xi_{I}$. The last equality holds because if $m\in Mon(f)$, it can be written as $m=m_{1}\cdot m_{2}\cdots m_{D_{2}}$ and for some $i\in[s]$ such that each sub-monomial $m_{j}\in X^{D_{1}}$, has a non-zero coefficient in $Q_{ij}$ and it can be processed by the substitution automaton independent of other sub-monomials. Since we sum over all possible computational paths for each monomial in $m\in Mon(f)$, the equality holds. Also suppose $m$ is computed by $\prod_{j\in[D_{2}]}Q_{i,j}$ for some $i\in[s]$ then by definition $m\in Mon(f)$. This completes the proof. ∎